if a beam is12"x12"x 23' long and weighs 3289lbs. 143bs per foot. how much weight is applied if the beam was to be lifted straight up from one end
1
Expert's answer
2020-12-06T17:31:45-0500

"W\frac{L}{2}=FL\\F=\frac{W}{2}=\frac{3289}{2}=1644.5\ lb_s"